OpenStudy (anonymous):

20x^5/15y^6 * 5y^4/6x^2 :)

OpenStudy (psymon):

Think you can cancel some things out? Thats what it looks like this whole problem is is the ability to reduce.

OpenStudy (anonymous):

But I thought you couldn't cancel things unless they had the same exponent?

OpenStudy (psymon):

20x^5/15y^6 * 5y^4/6x^2 \[\frac{ (20x^{5})(5y^{4}) }{ (15y^{6})(6x^{2}) }\]Thats really whats going on, you have all that stuff crammed together top and bottom as factors. You can surely cancel things out, x's with x's, y's with y's, just because theyre all being multiplied and on opposite sides of the fraction.

OpenStudy (anonymous):

so would it be like\[\frac{ (20)(20)(20)(20)(20)(5)(5)(5)(5) }{ (15)(15)(15)(15)(15)(15)(6)(6) } \] ?

OpenStudy (psymon):

More like this: \[\frac{ (20)(x)(x)(x)(x)(x)(5)(y)(y)(y)(y) }{ (15)(y)(y)(y)(y)(y)(y)(6)(x)(x) }\]

OpenStudy (anonymous):

\[\frac{ 20x^3(5) }{ 90 }\] ?

OpenStudy (psymon):

Everything looks right except there should be 2 y's on bottom. You can reduce the numbers further, too.

OpenStudy (anonymous):

10x^3/9y^2?

OpenStudy (psymon):

Looks fine to me :3

OpenStudy (anonymous):

YAY! thank you (again!) :D

OpenStudy (psymon):

Yeah, np :3